The range of the function f(x, y) =√81-9x²-9y² is [0, 8] . Say true or false.


1
Expert's answer
2022-01-24T16:45:27-0500
"f(x, y)=\sqrt{81-9x^2-9y^2}"

"x^2+y^2\geq0, x, y\in \R"

"-9x^2-9y^2\leq0"

"81-9x^2-9y^2\leq81"

We need that "81-9x^2-9y^2\geq 0"

Then


"0\leq\sqrt{81-9x^2-9y^2}\leq\sqrt{81}"

"0\leq\sqrt{81-9x^2-9y^2}\leq9"

The range of the function "f(x, y)" is "[0,9]."

Therefore the statement that the range of the function

"f(x, y)=\sqrt{81-9x^2-9y^2}"

is "[0, 8]" is False.